There was a major power outage on campus this morning, affecting two key server rooms. This resulted in the loss of most of the IT services on campus.
The majority of services were restored by 14:00. Notable exceptions being;
- Wi-Fi authentication
- Campus card payment services
- Sports Park web site
- Network link to Cedar Farm
- Some networking in Harry Pitt
- Lyle building power – affects all servers hosted in this server room
Some services are running slowly due to re-synchronisation of the server storage. This should complete over the weekend.
There is a backlog of e-mail (incoming and outgoing) which is being processed.
Please accept our apologies for the disruption this has caused.
By now you should be aware that some major changes are being introduced to IT support arrangements within the University. These changes form part of the Efficiency and Effectiveness Programme which is introducing new ways of working for us all.
This briefing information is intended for all staff and their managers. It summarises how these changes will affect you, and includes a quick guide on how to contact IT staff if you need help or assistance.
Please bring this to the attention of all who are likely to need IT support within your School or Department. Continue reading
Last Friday (15 Feb 2013) a phishing email was sent to a large number of staff at the University using a mocked up log-on page as below. Continue reading
IT Services would like to apologise for the problems with several of our services yesterday which will have caused disruption for staff and students.
Some services were noticeably slow early in the morning but by about 9.30 many key services (staff email, Blackboard, Agresso and collaborative shares) were unusable for many users. Part of the central file server infrastructure used by these services was rebooted and the Exchange mailboxes were moved to another file server which restored most of the services by about 11.30.
Unfortunately, Agresso took longer to restore and was not back until 14.20.
FMD P: drives were also partly affected and this service was rebooted yesterday evening.
ITS is continuing to treat this as a major incident and we are working with the supplier of our central file server on identifying the cause of the problem.
If you are still experiencing problems this morning, please inform ITS Help via x6262 or email@example.com
Wednesday 8 August 2012
All the computers in the S@il area on the first floor of the library will have no network connection from 730am until 430pm.
Thursday 9 August 2012
All the computers in the S@il area on the ground floor of the library will have no network connection from 730am until 1130am.
This is for essential maintenance to be carried out.
We would like to apologise for any inconvenience.